OpenStudy (anonymous):

Find the prime factorization of 50. Write the numbers with an "x" between them, like 2x3x7.

OpenStudy (anonymous):

50/2=25 25/5=5 5/5=1 2x5x5=50

OpenStudy (anonymous):

30m^2n isnt this 5x3m^2n

OpenStudy (anonymous):

m? n? there's no such thing

OpenStudy (anonymous):

thats what the worksheet has.